How do you solve this problem?

Find three consecutive whole numbers such that twice the sum of the two smallest numbers is 18 more than three times the largest number.
______ (smallest value)
______ (largest value)
1
Expert's answer
2011-02-17T05:36:46-0500
Denote these three numbers as x, x+1 and x+2.
The twice sum of two smallest numbers would be 2(x+(x+1)) = 18 + 3(x+2).
Let's find the x:
2x+2x+2 = 18 + 3x + 6
x = 22
x+1 = 23
x+2 = 24
smallest value is 22;
the largest value is 24